The initial concentration of a protein sample is 25 μg/μl. Calculate the volumes of the stock protein and water (diluent) necessary to prepare a 1:5 dilution in a final volume of 4 mL. What is the diluted concentration?

Volume of Protein:

Volume of Water:

Resulting Concentration:
Answer #1

I beleive the correct answer to be as follows:

1/5 = X/4,

X = 4/5 = 0.8mL.

So, you will need:

Volume of Protein: 800uL or 0.8mL

Volume of Water:3200uL or 3.2mL

Resulting Concentration: 5ug/ul.

feel free to leave a comment down below for any further query. good rating would be appreciated if you find my answer helpful. thank you.